three times the width of a certain rectangle exceeds twice its length by three inches and four times its length is twelve more than its perimeter.Find the dimensions of a rectangle.
Answer by edjones(7311)

A) 3w=2L+3
B) 4L=P+12
C) P=2L+2w
.
A)
w=2L/3+1
.
B)
P=4L-12
.
C)
4L-12=2L+2(2L/3+1)
4L-12=2L+4L/3+2
12L-36=6L+4L+6 multiply each side by 3 to eliminate the fraction.
2L=42
L=21"
.
B)
p=4*21-12
p=72"
.
A)
w=2*21/3+1
w=15"
.
Check:
C)
72=2*21+2*15 True
.
Ed